OpenStudy (anonymous):

Find the length of the hypotenuse of a right triangle with legs of 9 cm and 12 cm.

OpenStudy (aripotta):

Pythagorean theorem is a^2 + b^2 = c^2, so 9^2 + 12^2 = c^2

OpenStudy (anonymous):

thnx martha

OpenStudy (aripotta):

no problem :) do you know what to do from there?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

what ddo you do?

OpenStudy (aripotta):

what's 9 squared?

OpenStudy (aripotta):

...9 x 9 = ?

OpenStudy (anonymous):

81

OpenStudy (anonymous):

......

OpenStudy (aripotta):

and 12 x 12 = ?

OpenStudy (anonymous):

i know what to do

OpenStudy (aripotta):

ok then :l

OpenStudy (aripotta):

what's the answer?

OpenStudy (anonymous):

idk

OpenStudy (aripotta):

but you just said you knew what to do lol

OpenStudy (anonymous):

lol im using my big brothers acount im only 5

OpenStudy (anonymous):

so idk

OpenStudy (aripotta):

...but you told me you knew :/

OpenStudy (anonymous):

lol IM 5

OpenStudy (aripotta):

is this even your homework?

OpenStudy (anonymous):

how would i KNOW>!>!>!

OpenStudy (aripotta):

probably not

OpenStudy (aripotta):

bro. calm down. o-O

OpenStudy (anonymous):

aahhahahahaha

OpenStudy (anonymous):

i told you im five